# BED-LLM: Intelligent Information Gathering with LLMs and Bayesian Experimental Design

### v2

@@ 6.2 Preference Elicitation
Unlike 20 Questions, in which $\theta$ is a concrete entity and most reasonable questions have clear answers, many real-world information-gathering tasks involve more abstract targets and less predictable data generation. A key example is learning user preferences, where it may be difficult to explicitly define a concrete closed set of possible $\theta$, and it is also challenging for the LLM to develop appropriate uncertainty estimates. To study such a scenario, we evaluate BED-LLM on inferring users’ film preferences. Here the target we are trying to gather information about is somewhat abstract, and we have some flexibility in how we define $\theta$ in our joint model. Our chosen setup is to define $\theta$ to be a user profile, namely a paragraph of text describing the user’s film preferences, with our answerer model prompted to emulate a user with a given profile; see Appendix˜G for full details. We consider 200 different user profiles as the ground truth $\theta^{*}$, but as with 20 Questions this set is never given to the questioner. Because Split is not applicable as a baseline here (a deterministic likelihood assumption is clearly unreasonable), we benchmark with the similar Entropy approach instead. We also note that data–estimation setup is completely unviable here as well because of the large $\theta$ space.
At each turn $t\in(0,1,\ldots,5)$ we use the question-answer history as context for generating a list of ten film recommendations. This list is then rated in its fit to the user profile using an LLM-as-judge setup (Trivedi et al., 2024; Zhu et al., 2025). Specifically, the answerer scores each film on a scale of 1 to 5 (in 0.5 increments), based on how well the film aligns with $\theta^{*}$; this score is output together with a brief justification to increase reliability. The films’ scores are not included in $h_{t-1}$.
Our results in Figure˜2 show that, while Naive is often a strong baseline in this preference-elicitation scenario, BED-LLM is still able to provide a boost over both Naive and Entropy, producing higher-rated film recommendations. BED-LLM’s benefit is most clear in scenarios where the questioner belongs to a different model class to the answerer: here Naive’s performance is much less convincing.
@@ 7 Conclusion
In this work, we have shown how to effectively apply the framework of sequential Bayesian experimental design (BED) to the problem of interactive information gathering with LLMs. Specifically, we have introduced BED-LLM, which provides a specific, information–theoretic, sequential BED approach that makes a variety of carefully justified design choices in the joint-model factorization, belief updating, and EIG estimation. Particularly central to BED-LLM is the prior–likelihood pairing with filtering of hypotheses for consistency with the history. BED-LLM is notably the first work that uses both this prior–likelihood pairing without making a deterministic likelihood assumption that causes the EIG to simply to just marginal predictive entropy. Together, these innovations lead to substantial performance improvements compared to previous approaches. The results thus confirm that principled EIG–driven strategies can yield substantial gains for interactive, multi-turn, information gathering problems.
@@
Appendix A Algorithmic considerations
@@ A.1 Updating the likelihood
The success of BED-LLM hinges on our ability to update our joint distribution. As mentioned in §3, we choose not to update the likelihood model as more data is gathered, that is, our likelihood in the sequential setting will be $p_{\mathrm{LLM}}\left(y_{t};\left[\theta,x_{t}\right]\right)$ instead of $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},\theta,x_{t+1}])$. The main rationale of this choice is that for many problems our beliefs on $\theta$ capture all the required information to predict $y|x$, hence including the history is adding unnecessary context that could influence the LLM’s behavior in undesirable ways. However, it is important to note that $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},\theta,x_{t+1}])$ should be used instead for problems where $\theta$ will not capture all information from previous data, e.g. if $\theta$ is a binary value corresponding to whether we reject a null hypothesis, or is the answer to a particular other question of interest.
@@ A.2 Estimating EIG for each question
A.2.1 Predictive entropy is not a good approximation for EIG
@@
In practice, the expected likelihood entropy can and will vary across designs. This variability in the expected likelihood entropy can be crucial in selecting good designs. Here, we walk through a concrete example where predictive entropy might differ significantly from EIG.
Figure˜3 shows two candidate questions that could be asked to elicit film preference. Question 1 has high predictive entropy: in a randomly selected group of people, we would expect high variation in ice cream preference (regardless of the individual’s film preferences). However, since ice cream preference is unrelated to film preference, the answer would not help us narrow down our hypothesis space, and the EIG is zero.
@@ A.2.2 EIG estimator
One might be tempted to replace $\hat{p}(y_{t+1};[h_{t-1},x_{t+1}])$ with $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},x_{t+1}])$ in the EIG estimation in Equation˜3, as the two essentially offer alternative predictive distributions for the outcome. We also advise against this though, noting that it again provides no meaningful computational benefits (unless one also assumes a deterministic likelihood, but this would then mean we no longer consider $\theta$ at all). A key reason for avoiding this substitution is that it would mean we are no longer estimating a true EIG: the inconsistency between the likelihood and the marginal data distribution means there is no longer a joint model where we are minimising our expected uncertainty in $\theta$. We also find that the LLM process of sampling $\theta$ from $p_{f}(\theta;h_{t-1})$ followed by $y$ from $p_{\mathrm{LLM}}\left(y_{t};\left[\theta,x_{t}\right]\right)$ tends to give a better uncertainty over responses than sampling $y$ directly from $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},x_{t+1}])$.
@@ A.3 Prior construction and belief updating
In Section˜3.3, we argued that naive in-context updating is not sufficient for updating our beliefs: We fail to fully incorporate the information from the history $h_{t}$, and we often have overconfident distributions. The shortfalls of in-context learning in such settings have also previously be noted by, for example, (Liu et al., 2024b; Zhang et al., 2025; 2024). We posit two reasons why they likely struggle in such settings. First, the information from the different examples in the history are generally highly distinct in these information-gathering settings (indeed, this is part of our aim in adaptively design informative questions), making it harder for the LLM to appropriately reconcile all the provided information than in many other uses of in–context learning. Second, $\theta$ will often represent a user–specific variable that cannot easily be predicted from any data other than the user’s responses to questions: it has been argued that much of the success of in–context learning in LLMs is down to improving problem specification and linking the requested task to data it has seen in its training, rather than truly “learning” from the provided examples (Min et al., 2022; Kossen et al., 2024), but the history in our setting is rarely helpful for this due to its user–specific nature.
Appendix B Additional discussions on design choices
@@ B.1 An alternative view on the faithfulness of conditional distributions
Another way of viewing the distinction between the prior–likelihood and data–estimation constructions is in which of the EIG forms, Equation˜1 or Equation˜2, we center our reasoning. For a given joint model, the two are, of course, mathematically equivalent. However, they give us different ways of thinking about what it means to maximize the EIG: reducing entropy in $\theta$ from seeing $y$, or reducing entropy in $y$ from seeing $\theta$. This, in turn, gives us a way to reason about how appropriate our joint model is. When we choose to use one of $p\left(y_{t};\left[\theta,x_{t}\right]\right)$ or $p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$, we are centering our reasoning around the entropy of this quantity making sense, while allowing the other entropy in the other form to be implicitly defined from the resulting joint distribution; because the two forms are equivalent, we know that if our explicit form is suitable/unsuitable, the implicit form will be as well. If, for example, we directly fix the form of $p\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$ using our LLM’s predictive distribution, we are also directly relying on its expected entropy being a meaningful measure of design quality. If $\theta$ is high–dimensional and predominantly free–form, the resulting entropy produced by the LLM is unlikely to be meaningful and using the data–estimation pairing is unlikely to produce an effective strategy. However, if $y$ is instead quite constrained, the LLM can produce a meaningful entropy over it, and choosing a model based on the prior–likelihood pairing is likely to implicitly define a meaningful distribution, and thus entropy, on $\theta$. Conversely, if $\theta$ is constrained and $y$ is free form, the opposite will hold instead.
@@ B.3 Alignment between EIG and belief updating procedure
Our ultimate goal is to minimize uncertainty in $\theta$, as measured by its entropy. With this in mind, we can use the expected uncertainty reduction framework of Bickford Smith et al. (2025) to provide insights into how well our EIG formulation and belief updating procedures align.
To simplify discussions, for now we consider the setting where we choose a single question $x$ and obtain a response $y$. Following Bickford Smith et al. (2025), we can think of the “true” optimal design as selecting
$x^{*}_{\mathrm{true}}=\mathrm{argmin}_{x}\mathbb{E}_{p_{\mathrm{true}}(y;x)}\left[\mathrm{H}[p(\theta;x,y)]\right],$ (4)
@@
Note here that true optimal design has no direct dependency on our current beliefs about $\theta$; it only depends on $p_{\mathrm{true}}(y;x)$ and the hypothetical beliefs we produce for given observed data, $p(\theta;x,y)$. Thus, we can now see that our choice of joint model corresponds to different choices for approximating these quantities. Assuming that the LLM distribution is used directly for the conditional as per Section˜4, we thus have that:
  • The prior-likelihood pairing equates to the approximations $p_{\mathrm{true}}(y;x)\approx\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$ and $p(\theta;x,y)\approx p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])/\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$;
  • The data–estimation pairing equates to directly specifying a model for $p_{\mathrm{true}}(y;x)$ and then using the approximation $p(\theta;x,y)\approx p_{\mathrm{LLM}}(\theta;[x,y])$.
@@
The former of these considerations is difficult to control for as we simply do not know the true response distribution and it is hard to say which approach will thus estimate it best (though we can refer to the discussion in Section˜4 to determine which best matches our beliefs about the true response distribution). However, we do know upfront how we plan to derive our belief distribution on $\theta$ in practice, so we can use this to guide which joint model we formulate our EIG from. Namely, we observe that: a) using the prior–likelihood EIG pairing equates to assuming we will make a Bayesian update to our beliefs on $\theta$ using the likelihood $p_{\mathrm{LLM}}(y;[\theta,x])$; b) using the data–estimation EIG pairing equates to assuming we will make an in-context update to our beliefs on $\theta$, as we are treating $p(\theta;x,y)$ as $p_{\mathrm{LLM}}(\theta;[x,y])$.
Our preference between the pairings should therefore be guided in part by how we plan to update the model in practice. In particular, if we plan to make pure Bayesian updates, then the prior-likelihood formulation will tend to yield an EIG that is more faithful to our updating procedure, while if we only make simple in-context updates, the data–estimation formulation will tend to yield a more faithful EIG instead.
The update we use in practice, namely taking $p(\theta;x,y)=p_{f}(\theta;[x,y])$ as outlined in Section˜3.3, can be seen as being somewhere between the in-context and Bayesian updating: we initially sample from $p_{\mathrm{LLM}}(\theta;[x,y])$, but then perform filtering and other steps. The relative extent to which it resembles each will be problem–dependent and again be linked to how much we trust the LLM to capture uncertainty in the space of $\theta$ vs. $y$.
For the settings we consider, we expect $p_{f}(\theta;[x,y])$ to generally be better approximated by a Bayesian update than an in–context update, aligning with our decision to use the prior–likelihood formulation. The reasons for this are that a) the filtering often removes a large proportion of the generated samples, especially at later experiment turns, with $p_{\mathrm{LLM}}(\theta;h_{t-1})$ not fully incorporating information from the history; b) the maintaining of the set of one consistent hypotheses from one turn to the next encourages a more Bayesian behavior, with samples persisting unless contradicted by a new likelihood term; and c) the typical premature overconfidence of $p_{\mathrm{LLM}}(\theta;h_{t-1})$ to a small number of hypotheses means it is typically unrepresentative of our beliefs.
@@ Appendix C Extended related work
Several recent works have (explicitly or implicitly) looked at information gathering with LLMs. Most of these can be framed in a BED setting, with a deterministic likelihood (Piriyakulkij et al., 2023; Hu et al., 2024; Kobalczyk et al., 2025; Cooper et al., 2025), and can be seen as variants of our Split baseline. Piriyakulkij et al. (2023) use a deterministic 0/1 answer likelihood $p(a|x,q)$ via the LLM to prune items from a pre-enumerated finite set given a candidate question $q$. The question is selected by minimizing expected posterior entropy. They model user preferences with a binary ground truth, which would not be applicable in preference-elicitation scenarios with nebulous user profiles. Similarly,  Hu et al. (2024) use a deterministic likelihood to minimize entropy over a finite set $\Omega$ in a closed-world setting. Kobalczyk et al. (2025) target ambiguous task specifications in open-ended generation tasks by sampling a set of hypotheses (placing a uniform prior over them) and viewing each question as a deterministic partition over those samples, looking for questions that split the samples roughly evenly. Cooper et al. (2025) compute posterior entropy over a working set of top-k hypotheses (without filtering) through heuristic pruning.
@@
Algorithm 1 Data–Estimation Selection at Turn $t$
1:History $h_{t-1}$; candidate questions $\mathcal{X}_{\text{cand}}^{t}$; answer sets $\{\mathcal{Y}(x)\}$
2:Selected question $x_{t}^{\star}$
3:for each $x\in\mathcal{X}_{\text{cand}}^{t}$ do
4:  Obtain predictive answer distribution $p_{\text{LLM}}(y;x,h_{t-1})$ for all $y\in\mathcal{Y}(x)$
5:  for each $y\in\mathcal{Y}(x)$ do
6:   Compute entropy $H_{y}\leftarrow\mathrm{H}\!\left[p_{f}(\theta;h_{t-1}\cup(x,y))\right]$
7:  end for
8:  Compute $\text{EIG}(x)\leftarrow-\sum_{y\in\mathcal{Y}(x)}p_{\text{LLM}}(y;x,h_{t-1})\;H_{y}$
@@
10:Select $x_{t}^{\star}\leftarrow\arg\max_{x\in\mathcal{X}_{\text{cand}}^{t}}\text{EIG}(x)$
11:return $x_{t}^{\star}$

### v3

@@ 6.2 Preference elicitation
Unlike 20 Questions, in which $\theta$ is a concrete entity and most reasonable questions have clear answers, many real-world information-gathering tasks involve more abstract targets and less predictable data generation. A key example is learning user preferences, where it may be difficult to explicitly define a concrete closed set of possible $\theta$, or for the LLM to develop appropriate uncertainty estimates. To study such a scenario, we evaluate BED-LLM on inferring users’ film preferences. Here the target $\theta$ is somewhat abstract, and we have more flexibility in how we define it in our joint model. Our chosen setup is to define $\theta$ to be a user profile, namely a paragraph of text describing the user’s film tastes with our answerer model prompted to emulate a user with a given profile; see Appendix˜I for full details. We consider 200 ground-truth profiles, $\{\theta^{*}_{i}\}_{i=1}^{200}$, which, as with 20 Questions, are never revealed to the questioner. Because Split is not applicable as a baseline here (a deterministic likelihood assumption is clearly unreasonable), we benchmark with the similar Entropy approach instead. We also note that data–estimation setup is completely unviable here due to the large $\theta$ space.
At each turn, $t\in(0,1,\ldots,5)$, the questioner uses $h_{t-1}$ to generate ten film recommendations. This list is then rated in its fit to the user profile using an LLM-as-judge setup (Trivedi et al., 2024; Zhu et al., 2025). Specifically, the answerer scores each film on a scale of 1 to 5 (in 0.5 increments), based on how well the film aligns with $\theta^{*}$; this score is output together with a brief justification to increase reliability. The films’ scores are not included in $h_{t-1}$.
Our results in Figure˜3 show that, while Prompt-Only seems to be a stronger baseline in this preference-elicitation scenario than for 20 Questions, BED-LLM is still able to provide a boost over both it and Entropy, producing higher-rated film recommendations. BED-LLM’s benefit is most clear in scenarios where the questioner belongs to a different model class to the answerer.
@@ 7 Conclusion
In this work, we have shown how to effectively apply the framework of sequential Bayesian experimental design (BED) to the problem of interactive information gathering with LLMs. Specifically, we have introduced BED-LLM, which provides a specific, information-theoretic, sequential BED approach that makes a variety of carefully justified design choices in the joint-model factorization, belief updating, and EIG estimation. Particularly central to BED-LLM is the prior–likelihood pairing with filtering of hypotheses for consistency with the history. BED-LLM is notably the first work that uses both this prior–likelihood pairing without making a deterministic likelihood assumption that causes the EIG to simply to just marginal predictive entropy. Together, these innovations lead to substantial performance improvements compared to previous approaches. The results thus confirm that principled EIG-driven strategies can yield substantial gains for interactive, multi-turn information gathering problems.
Acknowledgments
DC is supported by the EPSRC CDT in Statistics and Machine Learning (EP/Y034813/1), an Oxford-Radcliffe Scholarship and an Oxford AI Fast Exploration grant. FBS is supported by the EPSRC Probabilistic AI Hub (EP/Y028783/1). TR is supported by the UK EPSRC grant EP/Y037200/1.
@@ Acknowledgments
Appendix A Discussion relating to Section˜3
@@ A.1 Updating the likelihood
The success of BED-LLM hinges on our ability to update our joint distribution. As mentioned in §3, we choose not to update the likelihood model as more data is gathered, that is, our likelihood in the sequential setting will be $p_{\mathrm{LLM}}\!\left(y_{t};\left[\theta,x_{t}\right]\right)$ instead of $p_{\mathrm{LLM}}(y_{t};[h_{t-1},\theta,x_{t}])$. The main rationale of this choice is that for many problems our beliefs on $\theta$ capture all the required information to predict $y|x$, hence including the history is adding unnecessary context that could influence the LLM’s behavior in undesirable ways. See Section˜F.3 for a results comparison of static and updated likelihoods on the 20 Questions game. However, it is important to note that $p_{\mathrm{LLM}}(y_{t};[h_{t-1},\theta,x_{t}])$ should be used instead for problems where $\theta$ will not capture all information from previous data, e.g. if $\theta$ is a binary value corresponding to whether we reject a null hypothesis, or is the answer to a particular other question of interest.
@@ A.2 Estimating EIG for each question
A.2.1 Predictive entropy is not a good approximation of EIG
@@
In practice, the expected likelihood entropy can and will vary across designs. This variability in the expected likelihood entropy can be crucial in selecting good designs. A concrete example helps highlight how predictive entropy can differ significantly from EIG. Figure˜4 shows two candidate questions that could be asked to elicit film preference. Question 1 has high predictive entropy: in a randomly selected group of people, we would expect high variation in ice cream preference (regardless of the individual’s film preferences). However, since ice cream preference is unrelated to film preference, the answer would not help us narrow down our hypothesis space, and the EIG is zero.
@@ A.2.2 EIG estimator
One might be tempted to replace $\hat{p}(y_{t+1};[h_{t-1},x_{t+1}])$ with $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},x_{t+1}])$ in the EIG estimation in Equation˜3, as the two essentially offer alternative predictive distributions for the outcome. We also advise against this, though, noting that it again provides no meaningful computational benefits (unless one also assumes a deterministic likelihood, but this would then mean we no longer consider $\theta$ at all). A key reason for avoiding this substitution is that it would mean we are no longer estimating a true EIG: the inconsistency between the likelihood and the marginal data distribution means there is no longer a joint model where we are minimising our expected uncertainty in $\theta$. We also find that the LLM process of sampling $\theta$ from $p_{f}(\theta;h_{t-1})$ followed by $y$ from $p_{\mathrm{LLM}}\!\left(y_{t};\left[\theta,x_{t}\right]\right)$ tends to give a better uncertainty over responses than sampling $y$ directly from $p_{\mathrm{LLM}}(y_{t+1};[h_{t-1},x_{t+1}])$.
@@ A.3 Prior construction and belief updating
In Section˜3.1, we argued that Prompt-Only in-context updating is not sufficient for updating our beliefs: We fail to fully incorporate the information from the history $h_{t}$, and we often have overconfident distributions. The shortfalls of in-context learning in such settings have also previously be noted by, for example, (Liu et al., 2024b; Zhang et al., 2025; 2024). We posit two reasons why they likely struggle in such settings. First, the information from the different examples in the history are generally highly distinct in these information-gathering settings (indeed, this is part of our aim in adaptively design informative questions), making it harder for the LLM to appropriately reconcile all the provided information than in many other uses of in–context learning. Second, $\theta$ will often represent a user–specific variable that cannot easily be predicted from any data other than the user’s responses to questions: it has been argued that much of the success of in–context learning in LLMs is down to improving problem specification and linking the requested task to data it has seen in its training, rather than truly “learning” from the provided examples (Min et al., 2022; Kossen et al., 2024), but the history in our setting is rarely helpful for this due to its user–specific nature.
Appendix B Discussion relating to Section˜4
@@ B.1 An alternative view on the faithfulness of conditional distributions
Another way of viewing the distinction between the prior–likelihood and data–estimation constructions is in which of the EIG forms, Equation˜1 or Equation˜2, we center our reasoning. For a given joint model, the two are, of course, mathematically equivalent. However, they give us different ways of thinking about what it means to maximize the EIG: reducing entropy in $\theta$ from seeing $y$, or reducing entropy in $y$ from seeing $\theta$. This, in turn, gives us a way to reason about how appropriate our joint model is. When we choose to use one of $p\!\left(y_{t};\left[\theta,x_{t}\right]\right)$ or $p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$, we are centering our reasoning around the entropy of this quantity making sense, while allowing the other entropy in the other form to be implicitly defined from the resulting joint distribution; because the two forms are equivalent, we know that if our explicit form is suitable/unsuitable, the implicit form will be as well. If, for example, we directly fix the form of $p\!\left(\theta;[h_{t-1},x_{t},y_{t}]\right)$ using our LLM’s predictive distribution, we are also directly relying on its expected entropy being a meaningful measure of design quality. If $\theta$ is high-dimensional and predominantly free-form, the resulting entropy produced by the LLM is unlikely to be meaningful and using the data–estimation pairing is unlikely to produce an effective strategy. However, if $y$ is instead quite constrained, the LLM can produce a meaningful entropy over it, and choosing a model based on the prior–likelihood pairing is likely to implicitly define a meaningful distribution, and thus entropy, on $\theta$. Conversely, if $\theta$ is constrained and $y$ is free form, the opposite will hold instead.
@@ B.3 Alignment between EIG and belief updating procedure
If our ultimate goal is to minimize uncertainty in $\theta$, as measured by its entropy, then we can use the expected uncertainty reduction framework of Bickford Smith et al. (2025) to provide insights into how well our EIG formulation and belief updating procedures align. To simplify discussions, for now we consider the setting where we choose a single question $x$ and obtain a response $y$. Following Bickford Smith et al. (2025), we can think of the “true” optimal design as selecting
$x^{*}_{\mathrm{true}}=\mathop{\mathrm{arg\,min}}_{x}\mathbb{E}_{p_{\mathrm{true}}(y;x)}\left[\mathrm{H}[p(\theta;x,y)]\right],$ (4)
@@
Note here that true optimal design has no direct dependency on our current beliefs about $\theta$; it only depends on $p_{\mathrm{true}}(y;x)$ and the hypothetical beliefs we produce for given observed data, $p(\theta;x,y)$. Thus, we can now see that our choice of joint model corresponds to different choices for approximating these quantities. Assuming that the LLM distribution is used directly for the conditional as per Section˜4, we thus have that
  • the prior–likelihood pairing corresponds to using the approximations $p_{\mathrm{true}}(y;x)\approx\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$ and $p(\theta;x,y)\approx p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])/\int p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])d\theta$; and
  • the data–estimation pairing corresponds to directly specifying a model for $p_{\mathrm{true}}(y;x)$ and then using the approximation $p(\theta;x,y)\approx p_{\mathrm{LLM}}(\theta;[x,y])$.
@@
The former of these considerations is difficult to control for as we simply do not know the true response distribution and it is hard to say which approach will thus estimate it best (though we can refer to the discussion in Section˜4 to determine which best matches our beliefs about the true response distribution). However, we do know upfront how we plan to derive our belief distribution on $\theta$ in practice, so we can use this to guide which joint model we formulate our EIG from. Namely, we observe that (a) using the prior–likelihood EIG pairing equates to assuming we will make a Bayesian update to our beliefs on $\theta$ using the likelihood $p_{\mathrm{LLM}}(y;[\theta,x])$, and (b) using the data–estimation EIG pairing equates to assuming we will make an in-context update to our beliefs on $\theta$, as we are treating $p(\theta;x,y)$ as $p_{\mathrm{LLM}}(\theta;[x,y])$.
Our preference between the pairings should therefore be guided in part by how we plan to update the model in practice. In particular, if we plan to make pure Bayesian updates, then the prior–likelihood formulation will tend to yield an EIG that is more faithful to our updating procedure, while if we only make simple in-context updates, the data–estimation formulation will tend to yield a more faithful EIG instead.
The update we use in practice, namely taking $p(\theta;x,y)=p_{f}(\theta;[x,y])$ as outlined in Section˜3.1, can be seen as being somewhere between the in-context and Bayesian updating: we initially sample from $p_{\mathrm{LLM}}(\theta;[x,y])$, but then perform filtering and other steps. The relative extent to which it resembles each will be problem–dependent and again be linked to how much we trust the LLM to capture uncertainty in the space of $\theta$ vs. $y$.
For the settings we consider, we expect $p_{f}(\theta;[x,y])$ to generally be better approximated by a Bayesian update than an in-context update, aligning with our decision to use the prior–likelihood formulation. The reasons for this are that a) the filtering often removes a large proportion of the generated samples, especially at later experiment turns, with $p_{\mathrm{LLM}}(\theta;h_{t-1})$ not fully incorporating information from the history; b) the maintaining of the set of one consistent hypotheses from one turn to the next encourages a more Bayesian behavior, with samples persisting unless contradicted by a new likelihood term; and c) the typical premature overconfidence of $p_{\mathrm{LLM}}(\theta;h_{t-1})$ to a small number of hypotheses means it is typically unrepresentative of our beliefs.
@@ Appendix C Extended related work
Several recent works have (explicitly or implicitly) looked at information gathering with LLMs. Most of these can be framed in a BED setting, with a deterministic likelihood (Piriyakulkij et al., 2023; Hu et al., 2024; Kobalczyk et al., 2025; Cooper et al., 2025), and can be seen as variants of our Split baseline. Piriyakulkij et al. (2023) use a deterministic 0/1 answer likelihood $p(a|x,q)$ via the LLM to prune items from a pre-enumerated finite set given a candidate question $q$. The question is selected by minimizing expected posterior entropy. They model user preferences with a binary ground truth, which would not be applicable in preference-elicitation scenarios with nebulous user profiles. Similarly,  Hu et al. (2024) use a deterministic likelihood to minimize entropy over a finite set $\Omega$ in a closed-world setting. Kobalczyk et al. (2025) target ambiguous task specifications in open-ended generation tasks by sampling a set of hypotheses (placing a uniform prior over them) and viewing each question as a deterministic partition over those samples, looking for questions that split the samples roughly evenly. Cooper et al. (2025) compute posterior entropy over a working set of top $k$ hypotheses (without filtering) through heuristic pruning.
@@
Algorithm 1 Data–Estimation sequential information gathering
1:LLM $p$, history $h_{0}=\varnothing$, budget $T$, num. hypothesis samples $M$
2:for $t=1,\ldots,T$ do
3:  Generate candidates: sample $\mathcal{X}_{\mathrm{cand}}=\{x_{t}^{(1)},\ldots,x_{t}^{(K)}\}$ from $p(x_{t};h_{t-1})$
4:  for each candidate $x_{t}\in\mathcal{X}_{\mathrm{cand}}$ do
5:   for each $y\in\mathcal{Y}$ do // enumerate answer options
6:     $w_{y}\leftarrow p(y;[h_{t-1},x_{t}])$ // answer prob. via LLM logits
7:     Sample $\{\theta_{m}^{(y)}\}_{m=1}^{M}\sim p(\theta;[h_{t-1},x_{t},y])$
8:     Estimate $\hat{H}_{y}\leftarrow\widehat{\mathrm{H}}\!\left[p(\theta;[h_{t-1},x_{t},y])\right]$ using $\{\theta_{m}^{(y)}\}$
@@
10:   $\hat{H}_{\mathrm{cond}}\leftarrow\sum_{y\in\mathcal{Y}}w_{y}\,\hat{H}_{y}$ // Term 1: marginal entropy of $\theta$ (design-dependent)
11:    Pool: $\{\theta_{m}\}_{m=1}^{M^{\prime}}\leftarrow\bigcup_{y\in\mathcal{Y}}\{\theta_{m}^{(y)}\}$ // samples from mixture
12:   for $m=1,\ldots,M^{\prime}$ do
13:     $\ell_{m}\leftarrow\log\sum_{y\in\mathcal{Y}}w_{y}\cdot p(\theta_{m};[h_{t-1},x_{t},y])$
14:   end for
15:    $\hat{H}_{\mathrm{marg}}\leftarrow-\frac{1}{M^{\prime}}\sum_{m=1}^{M^{\prime}}\ell_{m}$
16:   $\widehat{\mathrm{EIG}}(x_{t})\leftarrow\hat{H}_{\mathrm{marg}}-\hat{H}_{\mathrm{cond}}$
17:  end for
18:  Select: $x_{t}^{*}\leftarrow\arg\max_{x_{t}\in\mathcal{X}_{\mathrm{cand}}}\widehat{\mathrm{EIG}}(x_{t})$
19:  Ask and observe: pose $x_{t}^{*}$, receive $y_{t}$, set $h_{t}\leftarrow(h_{t-1},(x_{t}^{*},y_{t}))$
20:end for